x + 1=(square root of x + 43)
Answer by edjones(8007)

x+1=sqrt(x+43)
x^2+2x+1=x+43 Square each side.
x^2+x-42=0
(x+7)(x-6)=0
x=6 (x=-7 is an extraneous answer)
.
Ed